Home‎ > ‎Crew 159‎ > ‎


SOL: 5

12/18/2015 / Fri

HSO James A. F. Giammona


Second in Command: Chiara Cocchiara

Summary Title


Fossil Hunting



Crew 159 performed EVA4 with the primary goal of continuing to evaluate the new EVA suit in the field. The EVA4 crew also took this opportunity to hone their geology skills and look for fossils.


Morning briefings focused on 

1.       EVA4 agenda

2.       Checking in on how each crew member felt after a week of sim


Mission planning for the day


1.     0800 – Breakfast/ Plan for the day


2.     0905/0930 – Each crew member talked about how they felt the mission was going so far after a week.


3.     0930/1000 – XO Chiara Cocchiara described the plan for EVA4. After engineering checks, the EVA crew would pump water into the static tank and then head for a long walk northeast to look for interesting geology and fossils.


4.     1027/1225 – XO Chiara Cocchiara leads Crew Journalist Anna Ialeggio and Astroengineer Barak Stoltz on EVA4.


5.     1230/1300 – Lunchtime

A scrumptious meal of chili, pasta and bread was prepared by Commander Emmy Jewell.


6.     1300/1340 EVA4 debrief led by Commander Emmy Jewell. XO Cocchiara agrees with Susan Jewell’s assessment that the new EVA suit is more comfortable than the older suit. The crew was also excited by the fascinating terrain they encountered.


7.     1340/1800 – Crew has personal time to write reports and work on personal projects. Crew members also have researched their answers to the Open University students’ questions and will record video answers.



8.     1800/1845 – Dinner time and social time


9.     1900/2100 – CAPCOM time window


10.  2130/2230 Personal leisure time


11.  2300/0200 Astroengineer Barak Stoltz will work on calibrating the telescope and enabling remote operation. Tonight, he will work on integrating the CCD camera as last night the battery that rotates the dome was low on power.


Mission planning for the week 


1.     Crew Engineer Chiara Cocchiara will take the role of commander on Saturday at noon.


2.     Saturday 3 crew members depart (Susan Ip Jewell, Commander Nicholas Ip Jewell, and XO Emmy Helen Jewell). Crew 159 changes to 4 crew members on rotating Commander schedule.


3.     Crew 159 expect Joris Van Casteren, the journalist to visit us for the day on Saturday.


4.     Crew expects cameraman to arrive Monday or Tuesday.



Mission planning for the remaining time of the mission


1.     Remaining crew continues to work on personal projects


2.     Crew 159 will help Crew 160 travel from Grand Junction and assist in assembling their supplies.



1.     Another successful testing of the new suit during EVA4 led by XO Chiara Cocchiara. She was joined by Crew Journalist Anna Ialeggio and Astroengineer Barak Stoltz. After engineering checks, the EVA crew pumped water from the trailer to the static tank. Then they went on a long hike to practice mapping the geological features they encountered and to search for fossils.


2.     The crew researched and practiced delivering their answers for the Open University questions. They then recorded their responses on video.

3.     Astroengineer Stoltz worked on setting up remote control of the telescope and plans to install the CCD camera later tonight.


Look Ahead Plan:


1.         How are you going to follow up? 


1.     The EVA crew will continue to hone their geological mapping skills during EA5.


2.     We will continue to look at the chlorophyll samples with the Chi detector and will design a holder to integrate the detector with the rover.

2.       What is the plan for tomorrow?

1.       SOL 6 will see EVA5 which will involve more terrain exploration and practice identifying specific geomorphological features.

2.      We will also welcome the visiting journalist for the day.

Anomalies and any problems in work:  



Weather: A crisp, sunny day. Blue skies. Snow fully melted. No wind. 30-40 degrees F.

Crew Physical Status: Everyone remains in excellent health and high spirits.


EVA: where did you go?

After engineering check and transfer of water to static tank, EVA4 headed northeast and made a big loop back to the Hab for the next hour and a half.

Reports to be filed



EVA request

Daily Summary Report

HSO Report

Support Requested: